OpenStudy (anonymous):

Create an equation that results in at least one extraneous solution. Work through your equation, justify each step, and explain how the solution is extraneous.

OpenStudy (anonymous):

One type of equation that often gives extraneous solutions is one that contains absolute value. You could get a solution where the absolute value equals a negative number, etc.

OpenStudy (anonymous):

An example would be |x|=-1, absolute value is always positive, so any solution you come up with wouldn't work this will help you understand http://hotmath.com/hotmath_help/topics/extraneous-solutions.html

OpenStudy (anonymous):

Thanks @miryluslinares ! I've got a couple more questions for ya if you don't mind :) Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

One form is the square root of x to the fourth power: (x^4)^1/2 Another form is the fourth root of x to the eighth power: (x^8)^1/4 Another form is the eighth root of x to the sixteenth power: (x^16)^1/8 Each one can be simplified to x^2 by multiplying the exponent within the parenthesis with the external exponent.

OpenStudy (anonymous):

Wow you're good lol Thanks so much! Just one more :) Substitute your birthday into the equation \[\sqrt{x-y} + m = d\] where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be \[\sqrt{x-56} + 7 = 10\] Solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

im sorry i dont know how to do this one:(

OpenStudy (anonymous):

Oh, ok thanks anyway though!